# Multi-Agent Obstacle Avoidance using Velocity Obstacles and Control Barrier Functions

### v2

@@ Abstract
Velocity Obstacles (VO) methods form a paradigm for collision avoidance strategies among moving obstacles and agents. While VO methods perform well in simple multi-agent environments, they don’t guarantee safety and can show overly conservative behavior in common situations. In this paper, we propose to combine a VO-strategy for guidance with a CBF-approach for safety, which overcomes the overly conservative behavior of VOs and formally guarantees safety. We validate our method in a baseline comparison study, using 2nd order integrator and car-like dynamics. Results support that our method outperforms the baselines w.r.t. path smoothness, collision avoidance, and success rates.
@@ I Introduction
In this paper we propose a new combination VO and CBF for multi-agent collision avoidance. We use the classical CBF formulation of continuously solving an optimization problem to get a control input that is close to a desired one, while still guaranteeing safety. But instead of treating the VO as a constraint in this optimization, as [1], we move it into the objective function, as suggested in [2], while keeping a less restrictive CBF in the constraint to guarantee safety.
Multi-agent collision avoidance is a well studied problem with important applications in automated warehouses, autonomous driving and airborne drone delivery systems. The idea of VO was first proposed in [3], with a number of refinements in e.g. [4, 2, 1], and a nice recent survey in [5]. The key idea is to make the assumption that all other agents, at least temporarily, will keep a constant velocity, make sure your own velocity is such that no collisions occur, and keep repeating this as the world changes. This family of approaches have been shown to be very efficient in handling challenging scenarios with many agents involved. However, as was noted in [2], the approach might be overly conservative in some scenarios, where it leads to the conclusion that there are no admissible velocity choices left.
@@
Figure 1: An illustration of how VO terms in the strict CBF constraint can lead to very conservative behavior. The trajectories of two red agents aiming to reach a desired position are compared when subject to a strict (right) or a relaxed (left) Velocity Obstacle inequality constraint. The initial VO induced by the blue agent is illustrated as the light gray regions. The relaxed constraint allows the red agent on the left to pursue a more direct trajectory towards the destination, while the strict constraints makes it impossible for the red agent on the right to turn in front of the blue agent, since the turn at some point includes a velocity that would lead to a a collision if it was kept constant. Instead the red agent on the right has to pass behind the blue agent. In a scenario with more agents, similar concerns can lead to a very restricted set of options.
An efficient tool for guaranteeing safety in autonomous robotics is CBFs [7]. The central idea in CBFs is to formulate the safety constraint in terms of a scalar function, and then control the time derivative of that function, in a way similar to control Lyapunov functions [8], so that the system never reaches the un-safe region.
@@
The outline of the paper is as follows. In Section II we provide a brief background on CBFs and VO, and related work witin the area can be found in Section III. Then, in Section IV we present the proposed approach and the theoretical safety guarantees. The simulation results can be found in Section V, and conclusions can be found in Section VI.
@@ II Background
In this section we will describe the two approaches that we seek to combine. First we review the core result of CBFs, and then the VO approach. Finally, we briefly describe the motion models used in the examples.
@@ II-A Control Barrier Functions
CBFs have their roots in Lyapunov Theory [8], and a nice overview can be found in [7]. Let the system dynamics be control affine, i.e.,
@@
is non-empty for all $\mathbf{x}$, and if we choose controls $\mathbf{u}$ inside $K$, we are guaranteed to stay in the safe set $\mathbf{x}\in\mathcal{C}$, see [7]. Above, $\alpha$ is a so-called class $\mathcal{K}$ function, that is $\alpha:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+},\alpha(0)=0$ and $\alpha$ is strictly monotonic increasing, see Theorem 2 in [7]. Also note that sometimes the Lie-derivative notation is used, where
@@
If some desired control is given by $\mathbf{u}_{\mathrm{des}}=k(\mathbf{x})$, we can formulate the following optimization problem to find a control $\mathbf{u}$ that is close to $\mathbf{k}(\mathbf{x})$ while still inside $K$, thereby keeping the state inside the safe set $\mathcal{C}$, [7]
$\mathbf{u}(\mathbf{x})=\mbox{argmin}_{\mathbf{u}}$ $\quad\frac{1}{2}||\mathbf{u}-\mathbf{k}(\mathbf{x})||^{2}$ (4)
@@ II-B Velocity Obstacles
The idea of Velocity Obstacles (VO) was introduced in [3] and refined in e.g. [4, 2, 1]. The key idea is illustrated in Figure 2. If agent $B$ is moving with fixed velocity, we can study the avoidance problem from the perspective of agent $A$. If agent $A$ chooses the same velocity as agent $B$ (at the lower tip of the darker cone) the relative distance will not change. If a component is then added moving towards $B$, there will be a collision at some point. The higher up in the dark cone, the sooner the collision will happen (we might think of the lower tip as a collision at $t=\infty$). Note that the dark cone extends infinitely upwards (but this is hard to illustrate in a figure).
Formally the concept above gets a bit messy, but we can describe it as follows (allowing arbitrary shapes for $A$ and $B$ by using the $\oplus$-notation). First we define some set operations for $A,B\subset\mathbb{R}^{n}$ and $\mathbf{p},\mathbf{v}\in\mathbb{R}^{n}$.
$A\oplus B$ =$=\{\mathbf{a}+\mathbf{b}\mid\mathbf{a}\in A,\mathbf{b}\in B\},$ (6)
@@
Figure 2: Illustration of the key idea and notation in VO. If agent $A$ chooses a velocity outside the dark cone, it will not collide with agent $B$.
@@ II-C Vehicle motion models used in the simulations
where $p_{1},p_{2}$ is the position, $\theta$ represents the orientation of the car, $v$ is the speed, $L$ is the wheelbase distance, $\tan(\phi)$ is the steering control and $a$ is the acceleration control. Writing the above in the affine form $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+\mathbf{g}(\mathbf{x})\mathbf{u}$, with $\mathbf{x}=(p_{1},p_{2},\theta,v)$, and $\mathbf{u}=(\tan\phi,a)$ we get
@@ III Related Work
Collision avoidance in multi-agent systems is a huge field in itself, with subfields such as path planning [11], model predictive control (MPC) [12, 13] and artificial potential functions [14]. This paper is about merging the ideas of VO and CBFs so we will focus the description of related work in these two areas.
The idea of velocity obstacles was introduced in [3] and extended in a number of papers [15, 16, 4, 17], that have refined and developed the concept since. One identified issue with VO was the oscillation problem. If two agents are facing each other, then they might first select a velocity not leading to collision. But in the next timestep, they might conclude that no collision is imminent, given the current velocity of the other agent, and then (both) turn back to their initial velocities. This will create an oscillation, and works such as the Reciprocal Velocity Obstacle (RVO) [15] were proposed to resolve it. The RVO was later extended into the Hybrid RVO by [16] reducing the problem further, by making the velocity obstacles asymmetric, and built upon in [4, 17] where the problem is reduced to a linear programming problem that can be efficiently solved. In a paper on combining VO with MPC [13], the idea of including VO as constraints in the short horizon motion planning problem was explored.
@@
CBFs are a tool for guaranteeing safety in control system. Early works include [18], with an overview in [7], and an extension to higher order systems in [19]. In connection to avoidance of static obstacles, the connection between CBFs and artificial potential functions was explored in [20].
The work closest to ours within the CBF literature is [10] and [1], where the idea of combining CBFs and VO is proposed. However, in both papers the authors add the VO as a hard constraint, which can be overly restrictive, as observed in [2] and illustrated in Fig. 1. Thus our approach goes beyond [10, 1] in the sense that the VO constraint is relaxed, and thereby essentially moved into the objective of the optimization, while another (non-VO) collision avoidance component is added to the CBF to provide the safety guarantees.
@@ IV Method
Instead of using velocity obstacles to directly ensure safety, we employ them for guidance through a relaxed inequality constraint, and use a valid CBF to guarantee the safe behavior of agents, asuming that all agents seek to avoid collisions and that environments are non-degenerate. Altogether the optimization problem is formulated as
$\mathbf{u}_{i}=\mathop{\arg\min}\limits_{\mathbf{u}_{i}}$ $\quad k_{u}\lVert\mathbf{u}_{i}-\mathbf{u}_{\mathrm{ref},i}\rVert^{2}+k_{\mathrm{vo}}\sum_{j=1}^{n_{\mathrm{obs}}}w_{ij}\lambda_{ij}^{2}$ (16)
s.t. $\quad\dot{h}_{\mathrm{vo},ij}+\alpha_{\mathrm{vo}}(h_{\mathrm{vo},ij})\geq\lambda_{ij}$ (17)
$\quad\dot{h}_{c,ij}+\alpha_{c}(h_{c,ij})\geq 0$ (18)
$\quad\mathbf{u}\in\mathcal{U},$ (19)
where $\mathcal{U}=\{\mathbf{u}\in\mathbb{R}^{m}\mid\|\mathbf{u}\|\leq u_{\mathrm{max}}\}$ represents the valid input set. In this section we describe in detail the functions and variables composing the above optimization problem, starting with the objective function and then finishing with the constraints guaranteeing safety.
@@ IV-A Objective Function
The objective function for each agent is composed by a goal attractive component $J_{\mathrm{goal}}$ and a velocity-obstacle guidance component $J_{\mathrm{vo}}$. The goal attractive component is defined as the squared norm of the difference between the optimization variable to a reference input
@@
where the reference input $\mathbf{u}_{\mathrm{ref}}$ is given or computed by a known and state-dependent policy. In a scenario that is challenging with respect to the static obstacles, this would include planning a path that leads to the goal, and then tracking this path. The path tracking control would then be $\mathbf{u}_{\mathrm{ref}}$.
@@
$h_{\mathrm{vo},ij}(x,t)$ $=\mathbf{p}_{ij}^{T}\mathbf{v}_{ij}+\|\mathbf{p}_{ij}\|\|\mathbf{v}_{ij}\|\cos(\gamma_{ij})$ (21)
@@
Intuitively, the cone CBF constraint (22) keeps the relative vector $\mathbf{v}_{ji}$ outside of the Velocity Obstacle cone. Thus, satisfying (22) keeps the system safe, but it is overly constraining in many real scenarios, as pointed out in [2], and illustrated in Figure 1. As can be seen, this problem is significant in the case of a single obstacle, and gets even worse with multiple obstacles around, potentially making it very difficult to find a velocity outside of the joint set of Velocity Obstacles. The basic formulation used in [10] would result in the agent not finding any solution to the set of inequalities (other than the trivial 0 if its velocity is 0). To overcome this problem, we incorporate the VO as part of the objective by introducing an auxiliary (slack) variable $\lambda$ in Equation (22), giving
@@
$J_{i}=k_{u}J_{\mathrm{goal},i}+k_{\mathrm{vo}}\sum_{j=1}^{m}w_{ij}J_{\mathrm{vo},ij},$ (26)
with constant positive scaling factors $k_{u},k_{\mathrm{vo}}\in\mathbb{R}$. The weights are computed as a function of an approximate (linear extrapolation of positions) time-to-collision $T_{col,ij}$ between objects $i$ and $j$
$w_{ij}=\frac{1}{T_{col,ij}}.$ (27)
If objects $i$ and $j$ don’t collide according to the used model, then $T_{col,ij}\rightarrow\infty$ and we set $w_{ij}=0$. When using a linear model for extrapolation, the time-to-collision and therefore the weights are symmetric, i.e., $w_{ij}=w_{ji}$. The choice of an inverse time-to-collision for weighting encodes that an agent should dedicate more effort to avoid objects with which it would collide sooner if no further action is taken.
@@ IV-B Safety-Critical Constraint
The variable $d_{ij}$ represents the minimum distance between the $i$ and $j$ objects. We assume objects are convex, and therefore each of their distance functions is continuously differentiable. A valid CBF for double integrator systems with saturated input $\lVert\mathbf{u}\lVert\leq u_{\mathrm{max}}$ is
@@
$v_{ij}=\min\left(0,\mathbf{v}_{ij}^{T}\hat{\mathbf{p}}_{ij}\right),$ (29)
with the safety margin $\delta\geq 0$. The time derivative of the function (28) is linear on the acceleration. In the case where $\nu_{ij}<0$, we obtain the derivative
@@
and the corresponding linear inequality constraint
@@
where $\alpha_{c}$ is a class $\mathcal{K}$ function. The case where $\nu_{ij}=0$ arises only with a non-increasing distance between objects, and therefore the inequality constraint (31) is fulfilled for any acceleration.

### v3

@@ Abstract
Velocity Obstacles (VO) methods form a paradigm for collision avoidance strategies among moving obstacles and agents. While VO methods perform well in simple multi-agent environments, they do not guarantee safety and can show overly conservative behavior in common situations. In this paper, we propose to combine a VO strategy for guidance with a Control Barrier Function approach for safety, which overcomes the overly conservative behavior of VOs and formally guarantees safety. We validate our method in a baseline comparison study, using second-order integrator and car-like dynamics. Results support that our method outperforms the baselines with respect to path smoothness, collision avoidance, and success rates.
@@ I Introduction
In this paper we propose a combination of Velocity Obstacles (VO) and Control Barrier Functions (CBF) for multi-agent collision avoidance. We use the classical CBF formulation of continuously solving an optimization problem to get a control input that is close to a desired one, while still guaranteeing safety. Instead of treating the VO as a constraint in this optimization, as [1], we consider it in the objective function, as suggested in [2], while keeping a less restrictive CBF as a constraint to guarantee safety.
Multi-agent collision avoidance is a well-studied problem with important applications in automated warehouses, autonomous driving, and airborne drone delivery systems. The idea of VO was first proposed in [3], with a number of refinements in e.g. [4, 2, 1], and a nice recent survey in [5]. The key idea is to make the assumption that all other agents, at least temporarily, will keep a constant velocity, make sure your own velocity is such that no collisions occur, and keep repeating this as the world changes. This family of approaches have been shown to be very efficient in handling challenging scenarios with many agents involved. However, as was noted in [2], the approach might be overly conservative in some scenarios, where it leads to the conclusion that there are no admissible velocity choices left.
@@
Figure 1: An illustration showing that strict VO-based constraints can be overly conservative. Two red agents $A$ aim for a destination $A_{\mathrm{des}}$ while avoiding the obstacle $B$: the one with a relaxed constraint (left) follows a direct path, while the one with a strict constraint (right) cannot turn in front of the obstacle (whose initial VO is shown in light gray) as this would imply at some point selecting a velocity inside the VO. In more crowded scenarios, such strict constraints can severely limit available options.
An efficient tool for guaranteeing safety in autonomous robotics is CBFs [7]. The central idea in CBFs is to formulate the safety constraint in terms of a scalar function, and then control the time derivative of that function, in a way similar to control Lyapunov functions [8] so that the system never reaches the un-safe region.
@@
The outline of the paper is as follows. In Section II we provide a brief background on CBFs and VO, and related work within the area can be found in Section III. Then, in Section IV we present the proposed approach and the theoretical safety guarantees. The simulation results can be found in Section V, and conclusions are summarized in Section VI.
@@ II Background
In this section, we describe the two approaches that we seek to combine. First, we review the core result of CBFs, and then the VO approach. Finally, we briefly describe the motion models used in the examples.
@@ II-A Control Barrier Functions
CBFs have their roots in Lyapunov Theory [8], and an overview can be found in [7]. Let the system dynamics be control affine, i.e.,
@@
is non-empty for all $\mathbf{x}$, and if we choose controls $\mathbf{u}$ inside $K$, we are guaranteed to stay in the safe set $\mathbf{x}\in\mathcal{C}$, see [7]. The function $\alpha$ is a so-called class $\mathcal{K}$ function, that is ${\alpha:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+},}\,{\alpha(0)=0,}$ and $\alpha$ is strictly monotonic increasing, see Theorem 2 in [7]. Also note that sometimes the Lie-derivative notation is used, where
@@
If some desired control is given by $\mathbf{u}_{\mathrm{des}}=\mathbf{k}(\mathbf{x})$, we can formulate the following optimization problem to find a control $\mathbf{u}$ that is close to $\mathbf{k}(\mathbf{x})$ while still inside $K$, thereby keeping the state inside the safe set $\mathcal{C}$, [7]
$\mathbf{u}(\mathbf{x})=\operatorname*{argmin}_{\mathbf{u}}$ $\quad\frac{1}{2}||\mathbf{u}-\mathbf{k}(\mathbf{x})||^{2}$ (4)
@@ II-B Velocity Obstacles
The idea of Velocity Obstacles (VO) was introduced in [3] and refined in e.g. [4, 2, 1]. The key idea is illustrated in Figure 2. If agent $B$ is moving with fixed velocity, we can study the avoidance problem from the perspective of agent $A$. If agent $A$ chooses the same velocity as agent $B$ (at the lower tip of the darker cone) the relative distance will not change. If a component is then added moving towards $B$, there will be a collision at some point. The higher up in the dark cone, the sooner the collision will happen (we might think of the lower tip as a collision at $t=\infty$).
Formally, allowing arbitrary shapes for $A$ and $B$ by using the $\oplus$-notation, we can describe the concept above as follows. First, we define some set operations for $A,B\subset\mathbb{R}^{n}$ and $\mathbf{p},\mathbf{v}\in\mathbb{R}^{n}$
$A\oplus B$ =$=\{\mathbf{a}+\mathbf{b}\mid\mathbf{a}\in A,\mathbf{b}\in B\}$ (6)
@@
Figure 2: Illustration of the key idea and notation used in Velocity Obstacle methods. If agent $A$ chooses a velocity $\mathbf{v}_{A}$ outside the dark cone $VO_{B}^{A}(\mathbf{v}_{B})$, it will not collide with agent $B$.
@@ II-C Vehicle motion models used in the simulations
where $p_{1},p_{2}$ is the position, $\theta$ represents the orientation of the car, $v$ is the speed, $L$ is the wheelbase distance, $\tan\phi$ is the steering control, and $a$ is the acceleration control. Writing the above in the affine form ${\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+\mathbf{g}(\mathbf{x})\mathbf{u}}$, with ${\mathbf{x}=(p_{1},p_{2},\theta,v)}$, and ${\mathbf{u}=(\tan\phi,a)}$ we get
@@ III Related Work
Collision avoidance in multi-agent systems is a vast field in itself, with subfields such as path planning [11], model predictive control (MPC) [12, 13] and artificial potential functions [14]. This paper is about merging the ideas of VO and CBFs so we will focus on the description of related work in these two areas.
The idea of velocity obstacles was introduced in [3] and extended in a number of papers [15, 16, 4, 17], that have refined and developed the concept since. One identified issue with VO was the oscillation problem. If two agents are facing each other, then they might first select a velocity that avoids collision. But in the next timestep, they might conclude that no collision is imminent, given the current velocity of the other agent, and then (both) turn back to their initial velocities. This will create an oscillation and works such as the Reciprocal Velocity Obstacle (RVO) [15] were proposed to resolve it. The RVO was later extended into the Hybrid RVO by [16] reducing the problem further, by making the velocity obstacles asymmetric, and built upon in [4, 17] where the problem is reduced to a linear programming problem that can be efficiently solved. In a paper on combining VO with MPC [13], the idea of including VO as constraints in the short-horizon motion planning problem was explored.
@@
CBFs are a tool for guaranteeing safety in control systems. Early works include [18, 19], with an overview in [7], and an extension to higher order systems in [20]. In connection to the avoidance of static obstacles, the connection between CBFs and artificial potential functions was explored in [21].
The works closest to ours within the CBF literature are [10] and [1], where the idea of combining CBFs and VO is proposed. However, in both papers, the authors add the VO as a hard constraint, which can be overly restrictive, as observed in [2] and illustrated in Fig. 1. Thus our approach goes beyond [10, 1] in the sense that the VO constraint is relaxed, and thereby essentially moved into the objective of the optimization, while another (non-VO) collision avoidance component is added to the CBF to provide the safety guarantees.
@@ IV Method
Instead of using velocity obstacles to directly ensure safety, we employ them for guidance through a relaxed inequality constraint and use a valid CBF to guarantee safe behavior, under the assumption that all agents seek to avoid collisions. Here, agent $i$ denotes the agent for which we want to solve the optimization problem, and $j$ indexes the dynamic obstacles. Altogether, the optimization problem for the agent $i$ is formulated as
$\mathbf{u}_{i}=\operatorname*{argmin}_{\mathbf{u}_{i}}$ $\quad k_{u}\lVert\mathbf{u}_{i}-\mathbf{u}_{\mathrm{ref},i}\rVert^{2}+k_{\mathrm{vo}}\sum_{j=1}^{n_{\mathrm{obs}}}w_{ij}\lambda_{ij}^{2}$ (16)
s.t. $\quad\dot{h}_{\mathrm{vo},ij}+\alpha_{\mathrm{vo}}(h_{\mathrm{vo},ij})\geq\lambda_{ij},$ (17)
$\quad\dot{h}_{c,ij}+\alpha_{c}(h_{c,ij})\geq 0,$ (18)
$\quad\mathbf{u}_{i}\in\mathcal{U},$ (19)
where $\mathcal{U}=\{\mathbf{u}\in\mathbb{R}^{m}\mid\|\mathbf{u}\|\leq u_{\mathrm{max}}\}$ represents the valid input set. In this section we describe in detail the functions and variables composing the above optimization problem, starting with the objective function (16) and then finishing with the constraints (17)--(19) guaranteeing safety.
@@ IV-A Objective Function
The objective function for each agent is composed of a goal attractive component $J_{\mathrm{goal}}$ and a velocity-obstacle guidance component $J_{\mathrm{vo}}$. The goal attractive component is defined as the squared norm of the difference between the optimization variable to a reference input
@@
where the reference input $\mathbf{u}_{\mathrm{ref}}$ is given or computed by a known and state-dependent policy. In a scenario that is challenging with respect to the static obstacles, this would include planning a path that leads to the goal and then tracking this path. The path tracking control would then be $\mathbf{u}_{\mathrm{ref}}$.
@@
$h_{\mathrm{vo},ij}(\mathbf{x})$ $=\mathbf{p}_{ij}^{T}\mathbf{v}_{ij}+\|\mathbf{p}_{ij}\|\|\mathbf{v}_{ij}\|\cos(\gamma_{ij})$ (21)
@@
Intuitively, the cone CBF constraint (22) keeps the relative vector $\mathbf{v}_{ji}$ outside of the Velocity Obstacle cone. Thus, satisfying (22) keeps the system safe, but it is overly constraining in many real scenarios, as pointed out in [2], and illustrated in Figure 1. As can be seen, this problem is significant in the case of a single obstacle and gets even worse with multiple obstacles around, potentially making it very difficult to find a velocity outside of the joint set of Velocity Obstacles. The basic formulation used in [10] would result in the agent not finding any solution to the set of inequalities (other than the trivial 0 if its velocity is 0). To overcome this problem, we incorporate the VO as part of the objective by introducing an auxiliary (slack) variable $\lambda$ in Equation (22), giving
@@
$J_{i}=k_{u}J_{\mathrm{goal},i}+k_{\mathrm{vo}}\sum_{j=1}^{n_{\mathrm{obs}}}w_{ij}J_{\mathrm{vo},ij},$ (26)
with constant positive scaling factors $k_{u},k_{\mathrm{vo}}\in\mathbb{R}_{+}$. The weights are computed as a function of an approximate time-to-collision $T_{col,ij}$ (based on the linear extrapolation of positions) between objects $i$ and $j$
$w_{ij}=\frac{1}{T_{\mathrm{col},ij}}.$ (27)
If objects $i$ and $j$ do not collide according to the used model, then $T_{\mathrm{col},ij}\rightarrow\infty$ and we set $w_{ij}=0$. When using a linear model for extrapolation, the time-to-collision and therefore the weights are symmetric, i.e., $w_{ij}=w_{ji}$. The choice of an inverse time-to-collision for weighting encodes that an agent should dedicate more effort to avoid objects with which it would collide sooner if no further action is taken.
@@ IV-B Safety-Critical Constraint
The variable $d_{ij}$ represents the minimum distance between the $i$ and $j$ objects. We assume objects are convex, and therefore each of their distance functions is continuously differentiable. This assumption is not restrictive as continuously differentiable approximations of distance functions are readily available [24, 25, 26]. A valid CBF for double integrator systems with saturated input $\lVert\mathbf{u}\lVert\leq u_{\mathrm{max}}$ is
@@
$\nu_{ij}=\min\left(0,\mathbf{v}_{ij}^{T}\hat{\mathbf{p}}_{ij}\right),$ (29)
where $\delta\geq 0$ is a safety margin, $\hat{\mathbf{p}}_{ij}$ is the normalized vector between the agent and the closest point on object $j$, which makes $\nu_{ij}$ the relative velocity of the other object projected onto this direction with large negative $\nu_{ij}$ indicating a potential problem.
In the case where $\mathbf{v}_{ij}^{T}\hat{\mathbf{p}}_{ij}\leq 0$ we get
@@
i.e., time derivative of the function (28) is linear in the acceleration. If $\mathbf{v}_{ij}^{T}\hat{\mathbf{p}}_{ij}>0$ we get $\dot{h}_{c,ij}=\dot{d}_{ij}=\mathbf{v}_{ij}^{T}\hat{\mathbf{p}}_{ij}>0$. Either way, the corresponding linear inequality constraint to guarantee safety will be
@@
where $\alpha_{c}$ is a class $\mathcal{K}$ function. Note that if $\mathbf{v}_{ij}^{T}\hat{\mathbf{p}}_{ij}>0$, Equation (31) is independent of $\mathbf{u}_{ij}$ and always satisfied.